OpenStudy (anonymous):

Find the solutions 1/2x^2+2x+3=0

OpenStudy (anonymous):

It's basic algebra figure it out.

OpenStudy (anonymous):

I wouldn't be asking for help if I knew how.

OpenStudy (anonymous):

Read the text book...

OpenStudy (anonymous):

Why are you being rude?

OpenStudy (anonymous):

Is that \[\frac{ 1 }{ 2 } x^2 + 2x+3=0\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Ok well you can just use expansion in this equation. Have you all learned that?

OpenStudy (anonymous):

No, I haven't. If I have, I don't remember it.

OpenStudy (anonymous):

I don't really get it.

OpenStudy (anonymous):

When the equation in this condition\[ax^{2}+bx+c=0\]

OpenStudy (anonymous):

SerikMB is better at explaining

OpenStudy (anonymous):

you can apply the formula \[x=\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (anonymous):

Well thank you for trying Amber

OpenStudy (anonymous):

I just plug it in?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

here you have a=1/2 b=2 and c=3

OpenStudy (anonymous):

note the sign plus/minus, so you can have 2 solutions at quadratic equations

OpenStudy (anonymous):

-\[-2\sqrt{-2}/1\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yes, i have noticed that, in this case you have no particular REAL solution

OpenStudy (anonymous):

So what do I do? Is no solution the answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you, I appreciate the help.